Do Not Touch

Milo is currently attempting to render a 3D plot in Mathematica. It started computing this plot Saturday at 4pm (yes, that’s over 24 hours ago). Please, allow it to complete its computation.

Once it completes rendering the plot, the plot needs to be written out in PostScript. This is a simple process, but I would perfer to do it myself (if anything gets messed up, I’d rather not have someone else to blame).

N.B. Saving the Mathematica session is not a valid alternative, it would be VERY large and take comparable time to reload and rerender as it took to render additionally.

Preferably, when it completes (i.e. stops saying “Running...” and displays a 3D image somewhat like a cube), please let me know so I can come and deal with it. I suspect, it will take another hour or so to render the PostScript file after the 3D plot is generated.

I should be reachable at one of the following:

here (perphaps) Tech. Sq. x3-2657 Dorm Room x5-6387 AI andre@meyer.ai.mit.edu Athena andre@[charon,athena].mit.edu